1. Compute for the amount of charge (q) of a 15 V battery with -45 J potential energy
  2. A metal sphere has 5x10^-9C static charge. Compute the voltage if the radius is equal to 0.05m. (k=8.99 x 10^9 N.m^2/C^2
  3. Find the Total capacitance of the three given capacitors in series connection, given their individual capacitances, 4.0 μF, 15.0 μF.
  4. Find the net capacitance of the three given capacitors connected in parallel, given their individual capacitances, 2.0 μF, 5.0, μF, 8.0 μF.
  5. A parallel plate capacitor was placed 0.002 m apart. Compute for the capacitance if the area each plate is equal to 0.025 m^2. (ε0 = 8.85x10^-12 F/m)


with computations :) thank you


1
Expert's answer
2022-04-13T09:21:37-0400

1.

"q=\frac AU=-3~C."

2.

"U=\frac{kq}r=900~V."

3.

"C=\frac{C_1C_2}{C_1+C_2}=3.2~\mu F."

4.

"C=C_1+C_2+C_3=15~\mu F."

5.

"C=\frac{\varepsilon\varepsilon_0 S}d=0.11~nF."